# Assessing the replicability of RCTs in RWE emulations

### v1

@@ Abstract
Methods: We compare the statistical properties of the sceptical $p$-value and the two-trials rule. We illustrate the performance of the different methods using real-world evidence emulations of randomized, controlled trials (RCTs) conducted within the RCT DUPLICATE initiative.
@@ 1 Background
Randomized controlled trials (RCTs) represent the gold standard for proofing the efficacy of a new therapy (Byar et al.,, 1976; Abel and Koch,, 1999; Concato et al.,, 2000). To reduce bias and guarantee a high internal validity of the results, an RCT is standardized to the highest degree and has strict in- and exclusion criteria. However, older, multi-morbid patients, women (particularly those lactating, pregnant, or able to become pregnant) or non-consenting patients are often excluded from RCTs (Van Spall et al.,, 2007; Food and Drug Administration,, 2018) and as a result, a broad generalizability is sometimes missing (Schneeweiss and Avorn,, 2005; Eichler et al.,, 2011). Thus, some vulnerable patient groups are under-represented in RCTs, resulting in a lack of evidence for treatment decisions.
@@
Real-world evidence (RWE) plays an increasing role in research and for regulatory decision making, particularly due to an increasing number of potential data sources (US Food and Drug Administration,, 2023; Franklin et al.,, 2019; Orsini et al.,, 2020). RWE is given as the clinical evidence based on so-called real-world data (RWD), i.e. data outside of RCTs given as data relating to patient health status or data routinely collected from different sources (Franklin et al.,, 2021). With the increase in purely virtual data sources (such as "digital twins" or "Metaverse©"), the term "RWD" is more and more criticized, as all study data come from the "real" world, leading to an increased use of the term "routine practice data" (Institute for Quality and Efficiency in Health Care,, 2020). For the sake of clarity, however, the term RWD will be used as a synonym for routine practice data in the following and includes all type of data outside of RCTs.
@@
Studies based on RWD can be used for post-market analysis of treatments in actual real-world care to ensure patient safety (Schneeweiss and Avorn,, 2005; Eichler et al.,, 2011; Franklin et al.,, 2019), especially for research questions or patients groups that are difficult to address otherwise (Liu and Panagiotakos,, 2022; Franklin and Schneeweiss,, 2017; Franklin et al.,, 2019; Booth et al.,, 2019). In the literature known as efficacy-effectiveness gap, some studies reported a low performance of drugs and treatment in daily clinical practice (Bhandari et al.,, 2004; Eichler et al.,, 2011), but systematic analyses failed to proof a general significant difference between the effect estimate in observational studies and in RCTs (Concato et al.,, 2000; Golder et al.,, 2011; Kuss et al.,, 2011; Lonjon et al.,, 2014; Anglemyer et al.,, 2014). It is therefore of interest to determine if and when RCTs and well-designed RWE studies, closely aligned with the original RCT, reach the same conclusions. The RCT DUPLICATE initiative (Randomized, Controlled Trials Duplicated Using Prospective Longitudinal Insurance Claims: Applying Techniques of Epidemiology) consists of the emulation of 32 RCTs with RWD to check if and under which circumstances RWD studies explicitly designed to emulate RCTs are useful to draw the same causal conclusions (Franklin et al.,, 2021; Wang et al.,, 2023). A standardized and transparent study protocol for each RWE study was developed, with pre-defined patient selection and definition of primary analyses and study measures. Moreover, all emulation studies were a-priori registered before analyses were done (Franklin et al.,, 2021; Wang et al.,, 2023). All emulations were based on data of up to three different US claims data and propensity score matching with 120 pre-defined possible confounders was performed (Franklin et al.,, 2021; Wang et al.,, 2023).
@@
It is common to distinguish direct and conceptual replications (Nosek and Errington,, 2017). A direct replication is an attempt to reproduce a previously observed result based on new data, collected with a protocol as close as possible to the original study. Only the study sample size can be different and is often larger in replication studies to ensure sufficient power to confirm the original finding. In a conceptual replication, the aim is to show the robustness of a finding with a different study sample or study conditions. An emulation of an RCT is a conceptual replication, as it tries to confirm the result from an RCT with observational RWD (Wang et al.,, 2017). The possible problem of confounding through the lack of randomisation is usually addressed with statistical adjustments such as propensity score matching (Wang et al.,, 2023).
@@
Whether direct or conceptual, the evaluation of the replicability of two studies (e.g. RCT vs replication given by the RWE emulation) is not straightforward and different statistical methods are available (Held,, 2020; Errington et al.,, 2021). The regulatory agreement (also known as the two-trials rule) is a commonly used method to assess replicability of the two studies (Errington et al.,, 2021; Franklin et al.,, 2021). Regulatory agreement is fulfilled if the original effect and replication effect go in the same direction and are both significant. However, the pure assessment of the significance of the original trial and the replication study (two-trials rule – TTR) has the disadvantage that it does not directly take into account the effect sizes of the two studies (Held et al., 2022b, ).
@@
Another approach –  the so-called “sceptical” $p$-value –  was recently developed (Held,, 2020; Held et al., 2022a, ; Micheloud et al.,, 2023) and represents an alternative to the two-trials rule. It depends on the $p$-values of both studies and also on the variance ratio of both effect sizes. The original trial and the replication study are thus no longer interchangeable, a property which is particularly important with regards to the emulation of RCTs using RWD.
The aim of this paper is to investigate how the sceptical $p$-value performs in this context. To do so, the sceptical $p$-value will be used to evaluate the agreement of RWE emulation of RCTs given from the RCT DUPLICATE initiative (Franklin et al.,, 2021; Wang et al.,, 2023) and the results will be compared to the two-trials rule. First, we describe the data from the RCT DUPLICATE initiative in Section 2. Section 3 reviews the sceptical $p$-value and discusses Type-I error control and power for replication studies. In Section 4, the sceptical $p$-value is used to evaluate the replicability of emulations of RCTs given by the data set of the RCT DUPLICATE initiative (Franklin et al.,, 2021; Wang et al.,, 2023) and compared to the two-trials rule. The results are discussed in Section 5; especially in comparison with the conclusions from the original RCT DUPLICATE initiative.
@@ 2 Data source: RCT DUPLICATE initiative
For the study at hand, data of 32 RCTs and related RWE emulations from the RCT DUPLICATE initiative were available (Franklin et al.,, 2021; Wang et al.,, 2023). Detailed information about the RCT DUPLICATE initiative can be found in the literature (Franklin et al.,, 2020, 2021; Wang et al.,, 2023; Heyard et al.,, 2024). In short, published and ongoing RCTs were selected if the trials were relevant to regulatory decision making and were potentially replicable using RWD (Franklin et al.,, 2020, 2021). RCTs were considered to be potentially replicable if they fulfilled critical aspects concerning study protocol, primary outcome, inclusion and exclusion criteria. RCTs were chosen to get a mix of superiority and non-inferiority trials with large and small magnitudes of effect sizes and trials with active comparators as well as trials with active treatment and placebo added to active standard of care therapies (Franklin et al.,, 2020). All RCTs were required to be sufficiently large and well powered to guarantee reliable results.
@@
For all emulation studies, the data of up to three different US claims data-sources were used to emulate the related RCTs: Optum Clinformatics (from 01/2004 to 03/2019), IBM Market Scan (from 2003 to 2017) and a subset of Medicare Part A, B and D (2011-2017 including all patients with a diabetes or heart failure diagnosis, 2009-2017 including all patients with a dispensation for an oral anticoagulant). Information about other insurance holders are not available for this database. As a result, 13 of the 32 RCTs were not emulated on the Medicare dataset (TRITON-TIMI, PLATO, ISAR-REACT5, TRANSCEND, ON-TARGET, HORIZON-PIVOTAL, VERO, P04334, D5896, IMPACT, POET-COPD and INSPIRE). The development of all emulation studies was standardized and transparent, with prospectively defined patient selection, primary analyses and registration (Franklin et al.,, 2021; Wang et al.,, 2023). Within a pre-defined time period, all available patient information in the claims data were included to the studies, with inclusion and exclusion criteria as closely as possible adopted from original RCTs. Moreover, a 1:1 nearest-neighbor propensity score matching with caliper width of 0.01 to control for more than 120 possible confounders was performed for all replication studies. Confounders were selected a priori and measured during 6 months before drug initiation. Detailed information about confounder selection can be found in Franklin et al., (2021).
@@
For the analyses presented here, the pooled estimates from fixed effects meta-analysis of the three different data sources were used for all emulation studies. These pooled estimates have also been analyzed in Heyard et al., (2024) with focus on differences between RCT and RWD effect size estimates due to design and population differences. For the original trials with a non-inferiority design, the margins were retrieved from the original protocols. Only original trials with hazard ratios (HRs) as the primary effect size were included, resulting in the exclusion of the LEAD2 trial due to a continuous outcome. In addition, two further trials (ISAR-REACT5 and VERO trial) were excluded, because a meta-analysis was not performed due to significant heterogeneity between the estimates of the three US claims data detected by Cochran’s $Q$-test. Thus, $N=29$ pairs of original RCTs and related emulation studies (pooled data) were included for further analyses, $11$ with a superiority and $18$ with a non-inferiority design. The effect estimates of 28 trials were in the intended direction (HR < 1), out of which 26 were statistically significant. The two non-significant trials (INSPIRE and TRANSCEND) had one-sided $p$-values of $p=0.34$ and $p=0.10$, respectively. One trial (PRONOUNCE) had an effect estimate in the opposite direction (HR > 1), and is one of the two trials (with CAROLINA) for which the emulation was performed before the results of the RCT were made public. Hazard ratios with 95% confidence interval of all 29 analyzed trials from the RCT DUPLICATE initiative are presented in Fig. 1.
@@
Figure 1: Forest plot comparing RCT and RWE effect estimates (HR with 95% CI) from the RCT DUPLICATE initiative (Franklin et al.,, 2021; Wang et al.,, 2023). Individual non-inferiority margins were added for non-inferiority trials. Three trials were excluded due to a continuous outcome (LEAD2) and significant heterogeneity between the estimates of the three US claims data (ISAR-REACT5 and VERO).
@@ 3.1 The sceptical $p$-value
The lack of a standard approach to define replication success motivated Held, (2020) to develop a new method: the sceptical $p$-value, which combines two principles: ‘the Analysis of Credibility’ by Matthews, 2001a ; Matthews, 2001b and a prior-data conflict assessment proposed by Box, (1980).
Assume that the effect estimate from the RCT is in the anticipated direction ($\mbox{HR}<1$, so $\hat{\theta}_{\mbox{\scriptsize RCT}}=\log(\mbox{HR})<0$) and is significant at level $\alpha$, i.e. one-sided $p$-value $p_{\mbox{\scriptsize RCT}}<\alpha$. This result from the RCT is challenged by a sceptical prior that renders the significant finding no longer convincing (lower limit of the posterior credible interval fixed at zero). The corresponding variance $\tau^{2}$ of the sceptical prior decreases with increasing $p$-value $p_{\mbox{\scriptsize RCT}}$, so RCTs with a larger $p$-value will lead to a wider sceptical prior.
In the second step, the sceptical prior needs to be compared with the RWE effect estimate $\hat{\theta}_{\mbox{\scriptsize RWE}}$ with standard error $\sigma_{\mbox{\scriptsize RWE}}$. Held, (2020) proposed to assess prior-data conflict using the prior-predictive distribution of $\hat{\theta}_{\mbox{\scriptsize RWE}}$ (Spiegelhalter et al.,, 2004, Section 5.8). This results in the test statistic $t_{\mbox{\scriptsize Box}}={\hat{\theta}_{\mbox{\scriptsize RWE}}}/{\sqrt{\tau^{2}+\sigma_{\mbox{\scriptsize RWE}}^{2}}}$, which is compared against a standard normal $\operatorname{N}(0,1)$ distribution to obtain the tail probability $p_{\mbox{\scriptsize Box}}=\Phi(t_{\mbox{\scriptsize Box}})$ where $\Phi(.)$ denotes the standard normal cdf and smaller values of $p_{\mbox{\scriptsize Box}}$ indicate a larger conflict between the replication effect $\hat{\theta}_{\mbox{\scriptsize RWE}}$ and the sceptical prior. Replication success at level $\alpha$ is achieved if there is significant conflict between the prior and the RWE estimate, i.e. $p_{\mbox{\scriptsize Box}}\leq\alpha$.
The approach needs some adaptions for non-inferiority trials. If $\delta$ denotes the non-inferiority margin on the log hazard ratio scale, then the sceptical prior is now centered at $\delta$. We then obtain the modified test statistic $t_{\mbox{\scriptsize Box}}=({\hat{\theta}_{\mbox{\scriptsize RWE}}-\delta})/{\sqrt{\tau^{2}+\sigma_{\mbox{\scriptsize RWE}}^{2}}}$.

### v2

@@ Abstract
Methods: We review the statistical properties of the sceptical $p$-value and compare those to the two-trials rule. We extend the methodology to non-inferiority trials and describe how to invert the sceptical $p$-value to obtain confidence intervals. We illustrate the performance of the different methods using real-world evidence emulations of randomized, controlled trials (RCTs) conducted within the RCT DUPLICATE initiative.
@@ 1 Background
Randomized controlled trials (RCTs) represent the gold standard for proofing the efficacy of a new therapy (Byar et al., 1976; Abel and Koch, 1999; Concato et al., 2000). To reduce bias and guarantee a high internal validity of the results, an RCT is standardized to the highest degree and has strict in- and exclusion criteria. However, older, multi-morbid patients, women (particularly those lactating, pregnant, or able to become pregnant) or non-consenting patients are often excluded from RCTs (Van Spall et al., 2007; Food and Drug Administration, 2018) and as a result, a broad generalizability is sometimes missing (Schneeweiss and Avorn, 2005; Eichler et al., 2011). Thus, some vulnerable patient groups are under-represented in RCTs, resulting in a lack of evidence for treatment decisions.
@@
Real-world evidence (RWE) plays an increasing role in research and for regulatory decision making, particularly due to an increasing number of potential data sources (US Food and Drug Administration, 2023; Franklin et al., 2019; Orsini et al., 2020). RWE is given as the clinical evidence based on so-called real-world data (RWD), i.e. data outside of RCTs given as data relating to patient health status or data routinely collected from different sources (Franklin et al., 2021). With the increase in purely virtual data sources (such as "digital twins" or "Metaverse©"), the term "RWD" is more and more criticized, as all study data come from the "real" world, leading to an increased use of the term "routine practice data" (Institute for Quality and Efficiency in Health Care, 2020). For the sake of clarity, however, the term RWD will be used as a synonym for routine practice data in the following and includes all type of data outside of RCTs.
@@
Studies based on RWD can be used for post-market analysis of treatments in actual real-world care to ensure patient safety (Schneeweiss and Avorn, 2005; Eichler et al., 2011; Franklin et al., 2019), especially for research questions or patients groups that are difficult to address otherwise (Liu and Panagiotakos, 2022; Franklin and Schneeweiss, 2017; Franklin et al., 2019; Booth et al., 2019). In the literature known as efficacy-effectiveness gap, some studies reported a low performance of drugs and treatment in daily clinical practice (Bhandari et al., 2004; Eichler et al., 2011), but systematic analyses failed to proof a general significant difference between the effect estimate in observational studies and in RCTs (Concato et al., 2000; Golder et al., 2011; Kuss et al., 2011; Lonjon et al., 2014; Anglemyer et al., 2014). It is therefore of interest to determine if and when RCTs and well-designed RWE studies, closely aligned with the original RCT, reach the same conclusions. The RCT DUPLICATE initiative (Randomized, Controlled Trials Duplicated Using Prospective Longitudinal Insurance Claims: Applying Techniques of Epidemiology) consists of the emulation of 32 RCTs with RWD to check if and under which circumstances RWD studies explicitly designed to emulate RCTs are useful to draw the same causal conclusions (Franklin et al., 2021; Wang et al., 2023). A standardized and transparent study protocol for each RWE study was developed, with pre-defined patient selection and definition of primary analyses and study measures. Moreover, all emulation studies were a-priori registered before analyses were done (Franklin et al., 2021; Wang et al., 2023). All emulations were based on data of up to three different US claims data and propensity score matching with 120 pre-defined possible confounders was performed (Franklin et al., 2021; Wang et al., 2023).
@@
It is common to distinguish direct and conceptual replications (Nosek and Errington, 2017). A direct replication is an attempt to reproduce a previously observed result based on new data, collected with a protocol as close as possible to the original study. Only the study sample size can be different and is often larger in replication studies to ensure sufficient power to confirm the original finding. In a conceptual replication, the aim is to show the robustness of a finding with a different study sample or study conditions. An emulation of an RCT is a conceptual replication, as it tries to confirm the result from an RCT with observational RWD (Wang et al., 2017). The possible problem of confounding through the lack of randomisation is usually addressed with statistical adjustments such as propensity score matching (Wang et al., 2023).
@@
Whether direct or conceptual, the evaluation of the replicability of two studies (e.g. RCT vs replication given by the RWE emulation) is not straightforward and different statistical methods are available (Held, 2020; Errington et al., 2021). The regulatory agreement (also known as the two-trials rule, see e.g. Held, 2024) is a commonly used method to assess replicability of the two studies (Errington et al., 2021; Franklin et al., 2021). Regulatory agreement is fulfilled if the original effect and replication effect go in the same direction and are both significant. However, the pure assessment of the significance of the original trial and the replication study (two-trials rule – TTR) has the disadvantage that it does not directly take into account the effect sizes of the two studies (Held et al., 2022).
@@
Another approach –  the so-called “sceptical” $p$-value –  was recently developed (Held, 2020; Held et al., 2022; Micheloud et al., 2023) and represents an alternative to the two-trials rule. It depends on the $p$-values of both studies and also on the variance ratio of both effect sizes. The original trial and the replication study are thus no longer interchangeable, a property which is particularly important with regards to the emulation of RCTs using RWD.
The aim of this paper is to investigate how the sceptical $p$-value performs in this context. To do so, the sceptical $p$-value will be used to evaluate the agreement of RWE emulation of RCTs given from the RCT DUPLICATE initiative (Franklin et al., 2021; Wang et al., 2023) and the results will be compared to the two-trials rule. First, we describe the data from the RCT DUPLICATE initiative in Section 2. Section 3 reviews the sceptical $p$-value and extends it to non-inferiority studies. Type-I error control and power for replication studies are discussed and confidence intervals for the combined effect are derived. In Section 4, the sceptical $p$-value is used to evaluate the replicability of emulations of RCTs given by the data set of the RCT DUPLICATE initiative (Franklin et al., 2021; Wang et al., 2023) and compared to the two-trials rule. The results are discussed in Section 5; especially in comparison with the conclusions from the original RCT DUPLICATE initiative.
@@ 2 Data source: RCT DUPLICATE initiative
For the study at hand, data of 32 RCTs and related RWE emulations from the RCT DUPLICATE initiative were available (Franklin et al., 2021; Wang et al., 2023). Detailed information about the RCT DUPLICATE initiative can be found in the literature (Franklin et al., 2020, 2021; Wang et al., 2023; Heyard et al., 2024). In short, published and ongoing RCTs were selected if the trials were relevant to regulatory decision making and were potentially replicable using RWD (Franklin et al., 2020, 2021). RCTs were considered to be potentially replicable if they fulfilled critical aspects concerning study protocol, primary outcome, inclusion and exclusion criteria. RCTs were chosen to get a mix of superiority and non-inferiority trials with large and small magnitudes of effect sizes and trials with active comparators as well as trials with active treatment and placebo added to active standard of care therapies (Franklin et al., 2020). All RCTs were required to be sufficiently large and well powered to guarantee reliable results.
@@
For all emulation studies, the data of up to three different US claims data-sources were used to emulate the related RCTs: Optum Clinformatics (from 01/2004 to 03/2019), IBM Market Scan (from 2003 to 2017) and a subset of Medicare Part A, B and D (2011-2017 including all patients with a diabetes or heart failure diagnosis, 2009-2017 including all patients with a dispensation for an oral anticoagulant). Information about other insurance holders are not available for this database. As a result, 13 of the 32 RCTs were not emulated on the Medicare dataset (TRITON-TIMI, PLATO, ISAR-REACT5, TRANSCEND, ON-TARGET, HORIZON-PIVOTAL, VERO, P04334, D5896, IMPACT, POET-COPD and INSPIRE). The development of all emulation studies was standardized and transparent, with prospectively defined patient selection, primary analyses and registration (Franklin et al., 2021; Wang et al., 2023). Within a pre-defined time period, all available patient information in the claims data were included to the studies, with inclusion and exclusion criteria as closely as possible adopted from original RCTs. Moreover, a 1:1 nearest-neighbor propensity score matching with caliper width of 0.01 to control for more than 120 possible confounders was performed for all replication studies. Confounders were selected a priori and measured during 6 months before drug initiation. Detailed information about confounder selection can be found in Franklin et al. (2021).
@@
For the analyses presented here, the pooled estimates from fixed effects meta-analysis of up to three different data sources were used for all emulation studies. These pooled estimates have also been analyzed in Heyard et al. (2024) with focus on differences between RCT and RWD effect size estimates due to design and population differences. For the original trials with a non-inferiority design, the margins were retrieved from the original protocols. For the sake of clarity, only original trials with hazard ratios (HRs) as the primary effect size were included, resulting in the exclusion of the LEAD2 trial due to a continuous outcome. However, in general, the sceptical $p$-value can also be used for continuous or binary outcomes. In addition, two further trials (ISAR-REACT5 and VERO trial) were excluded, because a meta-analysis was not performed due to significant heterogeneity between the estimates of the three US claims data detected by Cochran’s $Q$-test. Thus, $N=29$ pairs of original RCTs and related emulation studies (pooled data) were included for further analyses, $11$ with a superiority and $18$ with a non-inferiority design. The effect estimates of 28 trials were in the intended direction (HR < 1), out of which 26 were statistically significant. The two non-significant trials (INSPIRE and TRANSCEND) had one-sided $p$-values of $p=0.34$ and $p=0.10$, respectively. One trial (PRONOUNCE) had an effect estimate in the opposite direction (HR > 1), and is one of the two trials (with CAROLINA) for which the emulation was performed before the results of the RCT were made public. Hazard ratios with 95% confidence interval of all 29 analyzed trials from the RCT DUPLICATE initiative are presented in Fig. 1 and appendix table B.1.
@@
Figure 1: Forest plot comparing RCT and RWE effect estimates (HR with 95% CI) from the RCT DUPLICATE initiative (Franklin et al., 2021; Wang et al., 2023). Individual non-inferiority margins were added for non-inferiority trials. Three trials were excluded due to a continuous outcome (LEAD2) and significant heterogeneity between the estimates of the three US claims data (ISAR-REACT5 and VERO).
@@ 3.1 The sceptical $p$-value
The lack of a standard approach to define replication success motivated Held (2020) to develop a new method: the sceptical $p$-value, which combines two principles: ‘the Analysis of Credibility’ by Matthews (2001a, b) and a prior-data conflict assessment proposed by Box (1980).
Assume that the effect estimate from the RCT is in the anticipated direction ($\mbox{HR}<1$, so $\hat{\theta}_{\mbox{\scriptsize RCT}}=\log(\mbox{HR})<0$) and is significant at level $\alpha$, i.e. one-sided $p$-value $p_{\mbox{\scriptsize RCT}}<\alpha$. This result from the RCT is challenged by a sceptical prior that renders the significant finding no longer convincing (lower limit of the posterior credible interval fixed at zero). The corresponding variance $\tau^{2}$ of the sceptical prior decreases with increasing (one-sided) $p$-value $p_{\mbox{\scriptsize RCT}}$, so RCTs with a larger $p$-value will lead to a wider sceptical prior. Note that the sceptical $p$-value for the alternative $\mbox{HR}<1$ is not available, if the effect estimate from the RCT is already in the opposite direction ($\mbox{HR}>1$), such as in the PRONOUNCE study.
In the second step, the sceptical prior needs to be compared with the RWE effect estimate $\hat{\theta}_{\mbox{\scriptsize RWE}}$ with standard error $\sigma_{\mbox{\scriptsize RWE}}$. Held (2020) proposed to assess prior-data conflict using the prior-predictive distribution of $\hat{\theta}_{\mbox{\scriptsize RWE}}$ (Spiegelhalter et al., 2004, Section 5.8). This results in the test statistic $t_{\mbox{\scriptsize Box}}={\hat{\theta}_{\mbox{\scriptsize RWE}}}/{\sqrt{\tau^{2}+\sigma_{\mbox{\scriptsize RWE}}^{2}}}$, which is compared against a standard normal $\operatorname{N}(0,1)$ distribution to obtain the tail probability $p_{\mbox{\scriptsize Box}}=\Phi(t_{\mbox{\scriptsize Box}})$ where $\Phi(.)$ denotes the standard normal cdf and smaller values of $p_{\mbox{\scriptsize Box}}$ indicate a larger conflict between the replication effect $\hat{\theta}_{\mbox{\scriptsize RWE}}$ and the sceptical prior. Replication success at level $\alpha$ is achieved if there is significant conflict between the prior and the RWE estimate, i.e. $p_{\mbox{\scriptsize Box}}\leq\alpha$.
The approach needs some adaptions for non-inferiority trials. If $\delta$ denotes the non-inferiority margin on the log hazard ratio scale, then the sceptical prior is now centered at $\delta$. We then obtain the modified test statistic
$t_{\mbox{\scriptsize Box}}=({\hat{\theta}_{\mbox{\scriptsize RWE}}-\delta})/{\sqrt{\tau^{2}+\sigma_{\mbox{\scriptsize RWE}}^{2}}}.$ (1)